# Twisted Superfields

### v8

@@ Introduction
The superspace gauge transformations of the model are built on the group U(3)$\times$×U(3). The field content of the model includes constructions built in N=1 superspace that are not N=1 superfields but are nonetheless called “twisted superfields” by way of analogy. Specifically, the model includes a real “twisted superfield”, an adjoint-representation chiral “twisted superfield”, and three flavors of fundamental and anti-fundamental chiral “twisted superfields”.
@@ 1 U(3) x U(3) Symmetries and Fields
The theory is constructed in N=1 superspace, extending four-dimensional spacetime by including four additional anticommuting coordinates $\theta_{\alpha},\bar{\theta}_{\dot{\alpha}}$. For reviews of superspace, see [6, 7, 8, 9, 10, 11, 12]; the notational conventions of [6] are used throughout. The gauge group of the model is U(3)$\times$×U(3), and it is described using 6$\times$6 matrices with the gauge fields in the 3$\times$3 diagonal blocks. The model includes a construction called a real “twisted superfield” that has the following attributes: field components in the 3$\times$3 diagonal blocks have an even number of $\theta_{\alpha},\bar{\theta}_{\dot{\alpha}}$ factors, while those in the 3$\times$3 off-diagonal blocks have an odd number of $\theta_{\alpha},\bar{\theta}_{\dot{\alpha}}$ factors. An adjoint-representation chiral “twisted superfield” has this same structure, and fundamental chiral “twisted superfields” have structures consistent with those.
@@
where each component field above is a U(3) matrix function of spacetime coordinates $x^{\mu}$. For example, $A_{1\mu}=A_{1\mu}^{A}\left(x\right)t^{A}$, where $t^{A}$ are 3$\times$3 U(3) matrices normalized by ${\rm tr}\left(t^{A}t^{B}\right)={\tfrac{1}{2}}\delta^{AB}$. Lower case letters are used to denote SU(3) adjoint indices $a,b\in\left\{1,2...8\right\}$. Upper case letters are used to denote U(3) adjoint indices $A,B\in\left\{0,1,2...8\right\}$ that include the Abelian matrix $t^{0}={\tfrac{1}{\sqrt{6}}}{\rm diag}\left(1,1,1\right)$. The $\theta_{\alpha}$ are 2-component anti-commuting Grassman coordinates, and $\bar{\theta}_{\dot{\alpha}}$ are their Hermitian conjugates. As a result of their $\theta_{\alpha},\bar{\theta}_{\dot{\alpha}}$ factors, the fields in the diagonal blocks of $V$ are bosons, while the fields in the off-diagonal blocks are fermions.
As mentioned previously, a theory built using the above real twisted superfield is not supersymmetric, since the fermion fields in the superfield are in a different representation of the U(3)$\times$×U(3) group than the boson fields. Despite not being supersymmetric, the real twisted superfield is assumed to transform as follows under a local “twisted supergauge transformation”:
@@
is a chiral “twisted superfield” whose component fields are U(3) matrix functions (e.g. $\alpha_{1}=\alpha_{1}^{A}t^{A}$) of $y^{\mu}=x^{\mu}+i\theta\sigma^{\mu}\bar{\theta}$. The twisted supergauge transformation of eq (1.2) maintains the boson-fermion structure of the real twisted superfield as well as its group structure. To the latter point, if the group was SU(3)$\times$×SU(3) instead of U(3)$\times$×U(3), the supergauge transformation would not be consistent, since a general supergauge transformation would generate terms in each block proportional to $t^{0}$. On the other hand, a U(3)$\times$×U(3) twisted supergauge transformation is consistent.
@@
$+i\sum_{k=1}^{\infty}\frac{B_{2k}}{\left(2k\right)!}\left[V,\left[V,...\left[V,\left(\Lambda^{{\dagger}}-\Lambda\right)\right]...\right]\right],$
@@
where ${\rm Tr}$ is the 6$\times$6 trace. The reason that this supergauge transformation is independent is because all of the commutators in eq (1.4) are proportional to some 6$\times$6 traceless matrix, so none of them can contribute to eq (1.6). Since by definition, the fermions of the real superfield are all in off-diagonal blocks, the field $V_{+}^{0}$ does not include any fermions, only bosons. Inside of $V$, the field $V_{+}^{0}$ is multiplied the 6$\times$6 matrix $T_{+}^{0}$ defined via
@@
This definition provides another way of saying that $T_{+}^{0}$ is ${\tfrac{1}{2\sqrt{3}}}$ of the 6$\times$6 unit matrix.
@@
In addition to the real superfield, the theory includes the following chiral twisted superfield in an adjoint representation of twisted U(3)$\times$×U(3):
@@
×$\times{\rm Tr}\Big{(}\Phi^{{\dagger}}e^{2gV^{\prime}}\Phi e^{-2gV^{\prime}}(1-\theta^{2}\bar{\theta}^{2}\sum_{m}m^{2}_{\Phi m}\sqrt{\tfrac{3}{2}}T_{m}^{0})\Big{)}$
@@
The $m^{2}_{\Phi m}$ terms are scalar mass terms. Despite the explicit group matrices $T_{m}^{0}$, the terms are supergauge invariant since the factor of $\theta^{2}\bar{\theta}^{2}$ limits gauge transformations to ones that remain within the same 3$\times$3 diagonal block. Just like the gaugino mass term (and like analogous soft supersymmetry breaking terms), the scalar mass terms break superspace translation invariance (which is not imposed in this model anyway), but do not break superspace gauge invariance.
In addition to adjoint-representation chiral fields, the theory also includes three flavors of 6-vector chiral twisted superfields in the fundamental and anti-fundamental representations of twisted U(3)$\times$×U(3):
@@
×$\hskip 11.38092pt\times\Big{(}\left(1-m^{2}_{mF}\theta^{2}\bar{\theta}^{2}\right)Q_{mF}^{{\dagger}}e^{2\left(gV^{\prime}+q_{mF}g_{+}V_{+}^{0}T_{+}^{0}\right)}Q_{mF}$
@@
×$\hskip 11.38092pt\times\left(\tilde{Q}_{mF}\left(m_{mFF^{\prime}}+\sqrt{2}\Gamma_{mFF^{\prime}}\Phi\right)Q_{mF^{\prime}}\right)$
@@
where $m\in\left\{1,2\right\}$ and $[mFF^{\prime}]$ means to only sum over combinations where $\tilde{Q}_{mF}$ and $Q_{mF^{\prime}}$ have the same $q_{mF}$ charge (and corresponding gauge transformations from eq (1.17)). The $m^{2}_{mF}$ and $\tilde{m}^{2}_{mF}$ terms generate additional mass terms for the fundamental and conjugate scalars.
@@
In [3], it was shown that a Wess-Zumino-like gauge is accessible for a theory with this kind of twisted supergauge invariance. After imposition of this Wess-Zumino gauge, the residual gauge invariance is just local spacetime gauge invariance. In the following, the fermions $\lambda$ and $\tilde{\lambda}$ will be referred to as “gauginos” despite the fact that they are in the $(3,\tilde{3})$ and $(\tilde{3},3)$ representations of the gauge group, rather than the adjoint representation.
Although the Abelian field $V_{+}^{0}$ has its own, independent gauge transformation, the second Abelian field $V_{-}^{0}$ (group structure $T_{-}^{0}$) does not decouple from the gauge transformations described above. That is why the same designation $g$ was used for both the $V_{-}^{0}$ and nonAbelian couplings above. Nonetheless, in the Wess-Zumino gauge, since $V_{-}^{0}$ is an Abelian field, it can accommodate different charges multiplying its coupling constant when acting on different chiral fields. This flexibility is used to make the charges of $V_{-}^{0}$ (labelled as $q^{\prime}_{mF}$)the same as those for $V_{+}^{0}$ defined in eq (1.18):
$q^{\prime}_{mF}=q_{mF}\,.$ (1.20)
The charge of $V_{-}^{0}$ is also $1$ when interacting with gauginos and the fermions in $\Phi$; it is only -2 when acting on fields with $mF=21$ or $1F$.
The classical theory has 50 parameters that can be adjusted classically: 2 coupling constants, 2 Abelian charges, 30 masses, 15 superpotential couplings, and a Fayet-Iliopoulos term. In the unification section of the paper, it is argued that both gauge couplings may be the same at the unification scale ($g_{+}=g$). It is also assumed that the following 8 parameters are zero classically:
$m_{23F}=m_{223}=\Gamma_{23F}=\Gamma_{223}=m_{23}=\tilde{m}_{23}=0$ (1.21)
@@ 2 Dynamical Symmetry Breaking
This section identifies a minimum of the scalar potential that breaks the gauge symmetry in stages from SU(3)$\times$×SU(3)$\times$×U(1)$\times$×U(1) down to SU(3)$\times$×U(1) and labels the fermions in the model based on their Standard Model symmetries.
@@
where $d_{\pm}^{0}={\tfrac{1}{\sqrt{2}}}\left(d_{1}^{0}\pm d_{2}^{0}\right)$, and lower-case ${\rm tr}$ defines a 3$\times$3 trace. By their equations of motion, the auxiliary fields are equal to linear or quadratic functions of the scalar fields. For example, the equations of motion for the d terms of $\mathcal{V}$ result in:
@@
$-d_{-}^{0}$ $=-{\frac{g}{\sqrt{12}}}\sum_{mF}q^{\prime}_{mF}\left(-1\right)^{m}\left(\phi_{mF}^{{\dagger}}\phi_{mF}-\tilde{\phi}_{mF}\tilde{\phi}_{mF}^{{\dagger}}\right)$
@@
The Abelian charges $q_{mF}$ and $q^{\prime}_{mF}$ for $V_{\pm}^{0}$ (including $d_{\pm}^{0}$) were defined in eqs (1.18) and (1.20). Just as for the $d$ auxiliary fields, the equations of motion can also be used to derive expressions for the $f$ auxiliary fields in terms of scalar fields.
@@
Due to eq (1.21) along with the assumption that $m_{21}$ and $m_{22}$ are small, it is assumed that the $m=2$ fundamental and conjugate scalars acquire vevs. Following precedent from Supersymmetric QCD (SQCD) [13, 12], the vevs in the $m=2$ sector are assumed to take the following form:
@@
where an overbar on a component of a scalar field (e.g. $\bar{\phi}_{2F}$ above) is used to denote the magnitude (real, positive) of the vev of that component. In eq (2.4), the index $n$ represents the SU(3) index of the 3-vectors $\tilde{\phi}_{2F}$ and $\phi_{2F}$. For example, writing out the SU(3) “color” components: $\left\langle\tilde{\phi}_{23}\right\rangle=i\left(0,0,\bar{\tilde{\phi}}_{23}\right)$. In other words, the vevs $\left\langle\tilde{\phi}_{2F}\right\rangle^{n}$ and $\left\langle\phi_{2F}\right\rangle_{n}$ form 3$\times$3 diagonal matrices in their flavor-“color” indices. The word “color” is being used here in order to make a connection with SQCD techniques, but in this model after symmetry breaking, the 3 “color” indices of the $m=2$ scalars will actually correspond to 2 isospin doublet indices and 1 singlet index. The phases of the vevs are chosen to simplify fermion mass matrices in the next section.
Unlike $m_{21}$ and $m_{22}$, it is assumed that the scalar masses $\tilde{m}_{21}$ and $\tilde{m}_{22}$ are large compared to $\xi_{+}$. To accommodate a nontrivial minimum in the presence of these masses, the following is assumed classically:
$\bar{\tilde{\phi}}_{21}=\bar{\tilde{\phi}}_{22}=0\hskip 14.22636pt\textrm{classically}.$ (2.5)
@@
$\left\langle\varphi_{2}\right\rangle$ $=\frac{\bar{\varphi}_{2}}{\sqrt{2}}\left(\begin{array}[]{ccc}{0}&{0}&{0}\\ {0}&{0}&{1}\\ {0}&{0}&{0}\end{array}\right)$
@@
In the next section, in light of the quantum vacuum moduli space, different assumptions will be made for both $\left\langle\varphi_{1}\right\rangle$ and $\left\langle\varphi_{2}\right\rangle$ that have ${\rm tr}\left\langle\varphi_{m}^{2}\right\rangle\neq 0$.
$2\left\langle\mathcal{V}\right\rangle$ $={\tfrac{1}{12}}g_{+}^{2}\left(\xi_{+}-2\bar{\phi}_{21}^{2}+\bar{\phi}_{22}^{2}+\Delta\bar{\phi}_{3}^{2}\right)^{2}$ (2.7)
$+{\tfrac{1}{6}}g^{2}\left(2\bar{\phi}_{21}^{2}-\Delta\bar{\phi}_{3}^{2}\right)^{2}+{\tfrac{1}{6}}g^{2}\bar{\phi}_{22}^{4}$
$+\tfrac{1}{4}g^{2}\left(\bar{\phi}_{22}^{2}-\Delta\bar{\phi}_{3}^{2}+\bar{\varphi}_{2}^{2}\right)^{2}$
$+2m_{21}^{2}\bar{\phi}_{21}^{2}+2m_{22}^{2}\bar{\phi}_{22}^{2}+\tilde{m}_{\Phi}^{2}\bar{\varphi}_{2}^{2}$
$\Delta\bar{\phi}_{3}^{2}$ $=\bar{\phi}_{23}^{2}-\bar{\tilde{\phi}}_{23}^{2}$
$\tilde{m}^{2}_{\Phi}$ $=m^{2}_{\Phi}+m^{2}_{\Phi 2}\,.$
It can be seen that if $\bar{\phi}_{23}^{2}=\bar{\tilde{\phi}}_{23}^{2}$ but all other vevs, masses and $\xi_{+}$ vanish, then an absolute minimum of the potential is attained: $\left\langle\mathcal{V}\right\rangle=0$. In this model, it is assumed that $\bar{\phi}_{23}$ and $\bar{\tilde{\phi}}_{23}$ are at a unification scale of $\sim 10^{16}$ GeV, but all other vevs, masses and $\xi_{+}$ in eq (2.7) are at the electroweak scale ($\sim 10^{2}$ GeV) or smaller. In that case, deviations from the absolute minimum value $\left\langle\mathcal{V}\right\rangle=0$ only at appear at the electroweak scale.
Well below the unification scale, the superpotential is more complicated since the couplings for different groups run differently. Nonetheless, by adjusting $\xi_{+}$, $\Delta\bar{\phi}_{3}^{2}$ and certain mass relations, classical minima can still be found where the electroweak scale is primarily defined by $\bar{\phi}_{21}$, and smaller scales are defined by $\bar{\phi}_{22}$ and $\bar{\varphi}_{2}$.
In particular, a minimum exists where three of the vevs are at the following scales:
$\begin{array}[]{|r|c|l|}\hline\cr\rm{vev}&\textrm{scale}&\mbox{Symmetry Breaking}\\ \hline\cr g\bar{\phi}_{23}&\sim 10^{16}\,\rm{GeV}&\mbox{Unification: SU(3)$\times$SU(3)$\times$U(1)$\times$U(1)$\to$SU(3)$\times$SU(2)$\times$U(1)$\times$U(1)}\\ g_{2}\bar{\phi}_{21}&\sim 10^{2}\,\rm{GeV}&\mbox{Electroweak: SU(3)$\times$SU(2)$\times$U(1)$\times$U(1)$\to$SU(3)$\times$U(1)$\times$U(1)}\\ g_{Z^{\prime}}\bar{\phi}_{22}&\sim 10^{-2}\,\rm{GeV}&\mbox{$Z^{\prime}$ acquires mass: SU(3)$\times$U(1)$\times$U(1)$\to$SU(3)$\times$U(1)}\\ \hline\cr\end{array}$ (2.8)
where $g_{2}$ is the Weak coupling and $g_{Z^{\prime}}$ is the $Z^{\prime}$ coupling that is driven to a very small value by an effective anomaly, as discussed in section 4. Since $\Delta\bar{\phi}_{3}^{2}$ is at the electroweak scale, but $g\bar{\tilde{\phi}}_{23}$ is at the unification scale, the following vev must also be at the unification scale:
$g\bar{\tilde{\phi}}_{23}\sim 10^{16}\rm{GeV}.$ (2.9)
@@
Via the Brout-Englert-Higgs mechanism, the scalar vevs $\bar{\phi}_{23}$ and $\bar{\tilde{\phi}}_{23}$ impart unification-scale masses to $A_{2\mu}^{4},A_{2\mu}^{5},A_{2\mu}^{6}$ and $A_{2\mu}^{7}$ as well as to one diagonal gauge boson. Consequently, the gauge symmetry is broken down to SU(3)$\times$×SU(2)$\times$×U(1)$\times$×U(1). To see the group structure of the remaining massless diagonal gauge fields, it is helpful to use the notation of eq (1.7) and re-expand the gauge fields into the linear combinations below:
@@
$q_{mF}g_{+}A_{+\mu}^{0}T_{+}^{0}+q^{\prime}_{mF}gA_{-\mu}^{0}T_{-}^{0}+gA_{2\mu}^{8}T_{2}^{8}=g_{Y}A_{\mu}^{Y}T_{mF}^{Y}+g_{Y^{\prime}}A_{\mu}^{Y^{\prime}}T_{mF}^{Y^{\prime}}+g_{U}A_{\mu}^{U}T_{mF}^{U},$ (2.11)
appearing in the action (eqs (1.18) and (1.20)), the coupling constants and group matrices must satisfy:
$\left(\begin{array}[]{c}{g_{Y}T_{mF}^{Y}}\\ {g_{Y^{\prime}}T_{mF}^{Y^{\prime}}}\\ {g_{U}T_{mF}^{U}}\end{array}\right)=\left(\begin{array}[]{ccc}{1}&{0}&{0}\\ {0}&{\cos\phi_{U}}&{\sin\phi_{U}}\\ {0}&{-\sin\phi_{U}}&{\cos\phi_{U}}\end{array}\right)\left(\begin{array}[]{ccc}{\cos\theta_{U}}&{0}&{\sin\theta_{U}}\\ {0}&{1}&{0}\\ {-\sin\theta_{U}}&{0}&{\cos\theta_{U}}\end{array}\right)\left(\begin{array}[]{c}{q_{mF}g_{+}T_{+}^{0}}\\ {q^{\prime}_{mF}gT_{-}^{0}}\\ {gT_{2}^{8}}\end{array}\right).$ (2.12)
This is just a generalization of a Weinberg angle rotation. A more complete generalization could involve a third angle mixing the two massless fields, but that is not needed here. The reason that the group matrices on the left have an $mF$ dependence is because eq (2.11) involves the Abelian charges $q_{mF}$ and $q^{\prime}_{mF}$.
@@
$\tan\phi_{U}$ $=\tfrac{1}{2}\cos\theta_{U}=\tfrac{1}{\sqrt{5}}.$

### v9

@@ Introduction
[…]
@@
In particular, the paper predicts the existence of a seventh quark, three heavy lepton families, a light $Z^{\prime}$ boson, a second Higgs boson, right-handed weak currents, and many additional scalars. Current experiments do not rule out these predictions; in fact data are better reproduced if they exist.
@@ 1 U(3) x U(3) Symmetries and Fields
[…]
@@
where each component field above is a U(3) matrix function of spacetime coordinates $x^{\mu}$. For example, $A_{1\mu}=A_{1\mu}^{A}\left(x\right)t^{A}$, where $t^{A}$ are $\times3$ U(3) matrices normalized by ${\rm tr}\left(t^{A}t^{B}\right)={\tfrac{1}{2}}\delta^{AB}$. Lower case letters are used to denote SU(3) adjoint indices $a,b\in\left\{1,2...8\right\}$. Upper case letters are used to denote U(3) adjoint indices $A,B\in\left\{0,1,2...8\right\}$ that include the Abelian matrix $t^{0}={\tfrac{1}{\sqrt{6}}}{\rm diag}\left(1,1,1\right)$. The $\theta_{\alpha}$ are 2-component anti-commuting Grassman coordinates, and $\bar{\theta}_{\dot{\alpha}}$ are their Hermitian conjugates. As a result of their $\theta_{\alpha},\bar{\theta}_{\dot{\alpha}}$ factors, the fields in the diagonal blocks of $V$ are bosons, while the fields in the off-diagonal blocks are fermions.
[…]
@@
[…]
@@
[…]
@@
where $\rm Tr$} is the $\times6$ trace. The reason that this supergauge transformation is independent is because all of the commutators in eq (1.4) are proportional to some $\times6$ traceless matrix, so none of them can contribute to eq (1.6). Since by definition, the fermions of the real superfield are all in off-diagonal blocks, the field $V_{+}^{0}$ does not include any fermions, only bosons. Inside of $V$, the field $V_{+}^{0}$ is multiplied the $\times6$ matrix $T_{+}^{0}$ defined via
@@
This definition provides another way of saying that $T_{+}^{0}$ is $\tfrac{1}{2\sqrt{3}}$} of the $\times6$ unit matrix.
@@
[…]
@@
$\times{\rm Tr}\Big(\Phi^{{\dagger}}e^{2gV^{\prime}}\Phi e^{-2gV^{\prime}}(1-\theta^{2}\bar{\theta}^{2}\sum_{m}m^{2}_{\Phi m}\sqrt{\tfrac{3}{2}}T_{m}^{0})\Big)$
@@
The $m^{2}_{\Phi m}$ terms are scalar mass terms. Despite the explicit group matrices $T_{m}^{0}$, the terms are supergauge invariant since the factor of $\theta^{2}\bar{\theta}^{2}$ limits gauge transformations to ones that remain within the same $\times3$ diagonal block. Just like the gaugino mass term (and like analogous soft supersymmetry breaking terms), the scalar mass terms break superspace translation invariance (which is not imposed in this model anyway), but do not break superspace gauge invariance.
[…]
@@
$\hskip 11.38092pt\times\Big(\left(1-m^{2}_{mF}\theta^{2}\bar{\theta}^{2}\right)Q_{mF}^{{\dagger}}e^{2\left(gV^{\prime}+q_{mF}g_{+}V_{+}^{0}T_{+}^{0}\right)}Q_{mF}$
@@
$\hskip 11.38092pt\times\left(\tilde{Q}_{mF}\left(m_{mFF^{\prime}}+\sqrt{2}\Gamma_{mFF^{\prime}}\Phi\right)Q_{mF^{\prime}}\right)$
@@
where $m\in\left\{1,2\right\}$ and $mFF^{\prime}]$ means to only sum over combinations where $\tilde{Q}_{mF}$ and $Q_{mF^{\prime}}$ have the same $q_{mF}$ charge (and corresponding gauge transformations from eq (1.17)). The $m^{2}_{mF}$ and $\tilde{m}^{2}_{mF}$ terms generate additional mass terms for the fundamental and conjugate scalars.
@@
In [3], it was shown that a Wess-Zumino-like gauge is accessible for a theory with this kind of twisted supergauge invariance. After imposition of this Wess-Zumino gauge, the residual gauge invariance is just local spacetime gauge invariance. In the following, the fermions $\lambda$ and $\tilde{\lambda}$ will be referred to as “gauginos” despite the fact that they are in the (3,3~)($\tilde{3}$) and $(\tilde{3},3)$ representations of the gauge group, rather than the adjoint representation.
Although the Abelian field $V_{+}^{0}$ has its own, independent gauge transformation, the second Abelian field $V_{-}^{0}$ (group structure $T_{-}^{0}$) does not decouple from the gauge transformations described above. That is why the same designation $g$ was used for both the $V_{-}^{0}$ and nonAbelian couplings above. Nonetheless, in the Wess-Zumino gauge, since $V_{-}^{0}$ is an Abelian field, it can accommodate different charges multiplying its coupling constant when acting on different chiral fields. This flexibility is used to make the charges of $V_{-}^{0}$ the same as those for $V_{+}^{0}$ defined in eq (1.18). In other words, the following replacement is made in that equation:
$q_{mF}g_{+}V_{+}^{0}T_{+}^{0}\to q_{mF}g_{+}V_{+}^{0}T_{+}^{0}+(q_{mF}-1)gV_{-}^{0}T_{-}^{0}\,.$ (1.20)
This change only affects fundamental and conjugate representation fields with with $mF=21$ or $1F$; it has no effect on fields in the real or adjoint twisted superfields. For interactions with gauge fields, this modification is equivalent to treating the $q_{mF}=-2$ fundamental or conjugate fields in each SU(3) sector as if they were in the antisymmetric representation of SU(3) for that sector.
The classical theory has 50 parameters that can be adjusted classically: 2 coupling constants, 2 Abelian charges, 30 masses, 15 superpotential couplings, and a Fayet-Iliopoulos term. In the unification section of the paper, it is argued that both gauge couplings may be the same at the unification scale ($g_{+}=g$). It is also assumed that the following 7 parameters are zero classically:
$m_{23F}=m_{223}=\Gamma_{23F}=m_{23}=\tilde{m}_{23}=0$ (1.21)
@@ 2 Dynamical Symmetry Breaking
[…]
@@
where $d_{\pm}^{0}={\tfrac{1}{\sqrt{2}}}\left(d_{1}^{0}\pm d_{2}^{0}\right)$, and lower-case $\rm tr$} defines a $\times3$ trace. By their equations of motion, the auxiliary fields are equal to linear or quadratic functions of the scalar fields. For example, the equations of motion for the d terms of $\mathcal{V}$ result in:
@@
$-d_{-}^{0}$ $=-{\frac{g}{\sqrt{12}}}\sum_{mF}q_{mF}\left(-1\right)^{m}\left(\phi_{mF}^{{\dagger}}\phi_{mF}-\tilde{\phi}_{mF}\tilde{\phi}_{mF}^{{\dagger}}\right)$
@@
The Abelian charges $q_{mF}$ for $V_{\pm}^{0}$ (including $d_{\pm}^{0}$) were defined in eq (1.18). Just as for the $d$ auxiliary fields, the equations of motion can also be used to derive expressions for the $f$ auxiliary fields in terms of scalar fields.
@@
Due to eq (1.21) along with an assumption that $m_{21}<\xi_{+}$, the $m=2$ fundamental and conjugate scalars may acquire vevs. Following precedent from Supersymmetric QCD (SQCD) [13, 12], the vevs in the $m=2$ sector are assumed to take the following form:
@@
where an overbar on a component of a scalar field (e.g. $\bar{\phi}_{2F}$ above) is used to denote the magnitude (real, positive) of the vev of that component. In eq (2.4), the index $n$ represents the SU(3) index of the 3-vectors $\tilde{\phi}_{2F}$ and $\phi_{2F}$. For example, writing out the SU(3) “color” components: $\left\langle\tilde{\phi}_{23}\right\rangle=i\left(0,0,\bar{\tilde{\phi}}_{23}\right)$. In other words, the vevs $\left\langle\tilde{\phi}_{2F}\right\rangle^{n}$ and $\left\langle\phi_{2F}\right\rangle_{n}$ form $\times3$ diagonal matrices in their flavor-“color” indices. The word “color” is being used here in order to make a connection with SQCD techniques, but in this model after symmetry breaking, the 3 “color” indices of the $m=2$ scalars will actually correspond to 2 isospin doublet indices and 1 singlet index. The phases of the vevs are chosen to simplify fermion mass matrices in the next section.
Unlike $m_{21}$, it is assumed that the scalar masses $m_{22}$, $\tilde{m}_{21}$ and $\tilde{m}_{22}$ are large compared to $\xi_{+}$. To accommodate a nontrivial minimum in the presence of these masses, the following is assumed classically:
$\bar{\phi}_{22}=\bar{\tilde{\phi}}_{21}=\bar{\tilde{\phi}}_{22}=0\hskip 14.22636pt\textrm{classically}.$ (2.5)
@@
$\left\langle\varphi_{2}\right\rangle$ $=-i\frac{\bar{\varphi}_{2}}{\sqrt{2}}\left(\begin{array}[]{ccc}{0}&{0}&{0}\\ {0}&{0}&{1}\\ {0}&{0}&{0}\end{array}\right)$
In the next section, in light of the quantum vacuum moduli space, a different assumption will be made for $\left\langle\varphi_{1}\right\rangle$, leading to ${\rm tr}\left\langle\varphi_{1}^{2}\right\rangle\neq 0$.
$2\left\langle\mathcal{V}\right\rangle$ $={\tfrac{1}{12}}g_{+}^{2}\left(\xi_{+}-2\bar{\phi}_{21}^{2}+\Delta\bar{\phi}_{3}^{2}\right)^{2}$ (2.7)
$+{\tfrac{1}{6}}g^{2}\left(2\bar{\phi}_{21}^{2}-\Delta\bar{\phi}_{3}^{2}\right)^{2}+\tfrac{1}{4}g^{2}\left(\bar{\varphi}_{2}^{2}-\Delta\bar{\phi}_{3}^{2}\right)^{2}$
$+2m_{21}^{2}\bar{\phi}_{21}^{2}+\tilde{m}_{\Phi}^{2}\bar{\varphi}_{2}^{2}$
$\Delta\bar{\phi}_{3}^{2}$ $=\bar{\phi}_{23}^{2}-\bar{\tilde{\phi}}_{23}^{2}\hskip 28.45274pt\tilde{m}^{2}_{\Phi}=m^{2}_{\Phi}+m^{2}_{\Phi 2}\,.$
It can be seen that if $\bar{\phi}_{23}^{2}=\bar{\tilde{\phi}}_{23}^{2}$ but all other vevs, masses and $\xi_{+}$ vanish, then an absolute minimum of the potential is attained: $\left\langle\mathcal{V}\right\rangle=0$. In this model, it is assumed that $\bar{\phi}_{23}$ and $\bar{\tilde{\phi}}_{23}$ are at a unification scale of $\sim 10^{14}$ GeV, but all other vevs, masses and $\xi_{+}$ in eq (2.7) are at the electroweak scale ($\sim 10^{2}$ GeV) or smaller. In that case, deviations from the absolute minimum value $\left\langle\mathcal{V}\right\rangle=0$ only at appear at the electroweak scale.
Well below the unification scale, the superpotential is more complicated since the couplings for different groups run differently. Nonetheless, by adjusting $\xi_{+}$, $\Delta\bar{\phi}_{3}^{2}$ and certain mass relations, classical minima can still be found where the electroweak scale is primarily defined by $\bar{\phi}_{21}$, and a smaller scale is defined by $\bar{\varphi}_{2}$.
In particular, a classical minimum exists with vevs at the following scales:
@@
where $g_{2}$ is the Weak coupling and $g_{Z^{\prime}}$ is the $Z^{\prime}$ coupling that is driven to a very small value by an effective anomaly, as discussed in section 4. Since $\Delta\bar{\phi}_{3}^{2}$ is at the electroweak scale, but $g\bar{\phi}_{23}$ is at the unification scale, the following vev must also be at the unification scale:
$g\bar{\tilde{\phi}}_{23}\sim 10^{14}\rm{GeV}.$ (2.9)
@@
[…]
@@
$q_{mF}g_{+}A_{+\mu}^{0}T_{+}^{0}+q_{mF}gA_{-\mu}^{0}T_{-}^{0}+gA_{2\mu}^{8}T_{2}^{8}=g_{Y}A_{\mu}^{Y}T_{mF}^{Y}+g_{Y^{\prime}}A_{\mu}^{Y^{\prime}}T_{mF}^{Y^{\prime}}+g_{U}A_{\mu}^{U}T_{mF}^{U},$ (2.11)
appearing in the action (eq (1.18)), the coupling constants and group matrices must satisfy:
$\left(\begin{array}[]{c}{g_{Y}T_{mF}^{Y}}\\ {g_{Y^{\prime}}T_{mF}^{Y^{\prime}}}\\ {g_{U}T_{mF}^{U}}\end{array}\right)=\left(\begin{array}[]{ccc}{1}&{0}&{0}\\ {0}&{\cos\phi_{U}}&{\sin\phi_{U}}\\ {0}&{-\sin\phi_{U}}&{\cos\phi_{U}}\end{array}\right)\left(\begin{array}[]{ccc}{\cos\theta_{U}}&{0}&{\sin\theta_{U}}\\ {0}&{1}&{0}\\ {-\sin\theta_{U}}&{0}&{\cos\theta_{U}}\end{array}\right)\left(\begin{array}[]{c}{q_{mF}g_{+}T_{+}^{0}}\\ {q_{mF}gT_{-}^{0}}\\ {gT_{2}^{8}}\end{array}\right).$ (2.12)
This is just a generalization of a Weinberg angle rotation. A more complete generalization could involve a third angle mixing the two massless fields, but that is not needed here. The reason that the group matrices on the left have an $mF$ dependence is because eq (2.11) involves the Abelian charges $q_{mF}$.
@@
$\tan\phi_{U}$ $=-\tfrac{1}{2}\cos\theta_{U}=-\tfrac{1}{\sqrt{5}}.$